[figure: 9.7 cm]

DETECTION OF TOROIDAL MOMENT IN Cr_{2}O_{3} AT THE SPIN - FLOP PHASE

Yu.F. Popov^{a}, A.M. Kadomtseva^{a}, A.K. Zvezdin^{b}, D.V. Belov^{a}, G.P. Vorob’ev^{a} and M.-M. Tehranchi^{b}

aDepartment of Physics, Moscow State University, 119899 Moscow, Russian Federation bInstitute of Physics, 117942 Moscow, Russian Federation

The magnetoelectric effect of Cr_{2}O_{3} single crystals had been investigated in a strong magnetic field up to 250 kOe. We have detected the existence of toroidal spin ordering in Cr_{2}O_{3} at the spin-flop phase, which is allowed by the symmetry of these crystals. This conclusion follows from the next experimental fact. The non-diagonal components of the magnetoelectric susceptibility tensor in Cr_{2}O_{3} have antisymmetric part which is dual to the toroidal moment. The toroidal moment first was observed experimentally in Ga_{2−x}Fe_{x}O_{3} crystals [1]. The toroidal moment in Cr_{2}O_{3}, in contrast to GaFeO_{3}, can be discovered only in the strong magnetic fields H||c when the spins cant to the ac plane and the non-diagonal components of the magnetoelectric tensor arise. Components of the electrical polarization P_{a} and P_{c} were measured in the magnetic field oriented in the ac plane at angle 5^{o} to c-axis. Comparison of the experimental curves with theoretical ones showed that non-diagonal components of magnetoelectric tensor (α_{13}, α_{31}) differ in sign. The significant asymmetry of the components α_{13} and α_{31} confirms that toroidal ordering exists in Cr_{2}O_{3}. This work is supported by RFBR (98-02-16848, 96-15-96429) and REM (97-0-7.3-157).

1. Yu.F. Popov, A.M. Kadomtseva, G.P. Vorob’ev at el., JETP 87, 146(1998)

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : Yu.F. Popov

Address for correspondence :

Department of Physics, Moscow State University, 119899 Moscow, Russian Federation

Email address : Popov@plms.phys.msu.su